OpenStudy (anonymous):

Simplify 5x + 2(x - 3) = -2(x - 1)

OpenStudy (anonymous):

5x + 2x -6 = -2x + 2 7x - 6 = -2x + 2 9x - 6 = 2 9x = 8 x = 8/9